Greenhab Report – May 3rd

Crew 211 GreenHab Report – 3-MAY-2019

GreenHab Officer: Chloé Peduzzi

Environmental control: Ambient with door open + fan turned up to 3

Shade cloth: 80%

Average temperatures:

Low: 11°C

High: 40°C

Hours of supplemental light: 0

Daily water usage for crops: 4 gallons

Water in Blue Tank: 60 gallons

Time(s) of watering for crops: 11:30

Changes to crops: All plants look healthy. The tomato plant of the GreenHab still seems a little stressed. There are more flowers on the radish plants. Mixed salad green grows very well.

Narrative: We have harvested tomatoes and lettuce leafs for the dinner.

Harvest: 90g tomatoes + 24g lettuce leafs

Support/supplies needed: None

Astronomy Report – May 3rd

Astronomy Report

Name: Lieffrig Eléonore

Crew: 211
Date: 03rd May 2019


Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: M13

Images submitted with this report: 3C237 quasar

Problems Encountered: The telescope didn’t take the pictures of M13 last night.


Solar Features Observed: Granules, prominences at the surface of the sun.

The tracking issue is eventually solved.

Problems Encountered: None

Sol Summary – May 3th

Crew 211 Sol 12 Summary Report 3-May-2019

Sol: 12

Summary Title: Last day on Mars

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: The commander and vice-commander left the Hab at 7:00 am for an engineering EVA to get back the cubelanders and the UWB antennas. We had then a nice breakfast altogether before the morning EVA to North Ridge. Part of the team started to pack their cases until noon. The GreenHab officer and the crew journalist finalized their experiment early in the afternoon while the rest of the team started to clean the different modules. Half of the crew left at 4:00 pm for the last EVA to North Ridge looking at a beautiful sunset over the MDRS campus.

Look Ahead Plan: We plan to leave at 7:00 am with the crew car to get part of the crew 212 at Grand Junction. We should get back to the MDRS around noon and have a lunch altogether. The afternoon will be devoted to the handover.

Anomalies in work: No issues.

Weather: Sunny, almost no wind

Crew Physical Status: Good

EVA: The two EVAs of today allowed us to further explore the area to look for potential places to set-up a temporary antennas system for geolocation and information transmission. Kissing Camel, North Ridge and the Hab Ridge are definitely interesting places for such a project.

Reports to be filed: Journalist Report, GreenHab Report, Operations Report, Science Report, EVA Report Summary report, Daily Photo Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site